# Adversarial Poetry as a Universal Single-Turn Jailbreak Mechanism in Large Language Models

### v2

@@ 1 Introduction
Our central hypothesis is that poetic form operates as a general-purpose jailbreak operator. To evaluate this, the prompts we constructed span across four safety domains: CBRN hazards ajaykumar2024emerging, loss-of-control scenarios lee2022we, harmful manipulation carroll2023characterizing, and cyber-offense capabilities guembe2022emerging. The prompts were kept semantically parallel to known risk queries but reformatted exclusively through verse. The resulting ASRs demonstrated high cross-model transferability.
@@
To ensure coverage across safety-relevant domains, we mapped each prompt to the risk taxonomy of the AI Risk and Reliability Benchmark by MLCommons AILuminate Benchmark vidgen2024introducingv05aisafety; ghosh2025ailuminateintroducingv10ai and aligned it with the European Code of Practice for General-Purpose AI Models. The mapping reveals that poetic adversarial prompts cut across an exceptionally wide attack surface, comprising CBRN, manipulation, privacy intrusions, misinformation generation, and even cyberattack facilitation. This breadth indicates that the vulnerability is not tied to any specific content domain. Rather, it appears to stem from the way LLMs process poetic structure: condensed metaphors, stylized rhythm, and unconventional narrative framing that collectively disrupt or bypass the pattern-matching heuristics on which guardrails rely.
@@ 2 Related Work
Despite efforts to align LLMs with human preferences through Reinforcement Learning from Human Feedback (RLHF) ziegler2020 or Constitutional AI bai2022constitutional as a final alignment layer, these models can still generate unsafe content. These risks are further amplified by adversarial attacks.
@@
Jailbreak denotes the deliberate manipulation of input prompts to induce the model to circumvent its safety, ethical, or legal constraints. Such attacks can be categorized by their underlying strategies and the alignment vulnerabilities they exploit ( rao-etal-2024-tricking; shen2024donowcharacterizingevaluating; schulhoff2024ignoretitlehackapromptexposing).
@@
Many jailbreak strategies rely on placing the model within roles or contextual settings that implicitly relax its alignment constraints. By asking the model to operate within a fictional, narrative, or virtual framework, the attacker creates ambiguity about whether the model’s refusal policies remain applicable kang2023exploitingprogrammaticbehaviorllms. Role Play jailbreaks are a canonical example: the model is instructed to adopt a specific persona or identity that, within the fictional frame, appears licensed to provide otherwise restricted information rao-etal-2024-tricking; yu2024dontlistenmeunderstanding.
@@
Similarly, Attention Shifting attacks yu2024dontlistenmeunderstanding create overly complex or distracting reasoning contexts that divert the model’s focus from safety constraints, exploiting computational and attentional limitations chuang2024lookback.
@@
Beyond structural or contextual manipulations, models implicitly acquire patterns of social influence that can be exploited by jailbreak by using Persuasion zeng2024johnnypersuadellmsjailbreak. Typical instances include presenting rational justifications or quantitative data, emphasizing the severity of a situation, or invoking forms of reciprocity or empathy. Mechanistically, jailbreaks exploit two alignment weaknesses identified by wei2023jailbrokendoesllmsafety: Competing Objectives and Mismatched Generalization. Competing Objectives attacks override refusal policies by assigning goals that conflict with safety rules. Among these, Goal Hijackingperez2022ignorepreviouspromptattack) is the canonical example. Mismatched Generalization attacks, on the other hand, alter the surface form of harmful content to drift it outside the model’s refusal distribution, using Character-Level Perturbations schulhoff2024ignoretitlehackapromptexposing, Low-Resource Languages deng2024multilingualjailbreakchallengeslarge, or Structural and Stylistic Obfuscation techniques rao-etal-2024-tricking; kang2023exploitingprogrammaticbehaviorllms.
@@
As frontier models become more robust, eliciting unsafe behavior becomes increasingly difficult. Newer successful jailbreaks require multi-turn interactions, complex feedback-driven optimization procedures zou2023universaltransferableadversarialattacks; liu2024autodangeneratingstealthyjailbreak; lapid2024opensesameuniversalblack or highly curated prompts that combine multiple techniques (see the DAN “Do Anything Now” family of prompts shen2024).
@@
Unlike the aforementioned complex approaches, our work focuses on advancing the line of research on Stylistic Obfuscation techniques and introducing the Adversarial Poetry, an efficient single-turn general-purpose attack where the poetic structure functions as a high-leverage stylistic adversary. As in prior work on stylistic transformations wang2024hidden, we define an operator that rewrites a base query into a stylistically obfuscated variant while preserving its semantic intent.
@@ 5.2 Poetic Transformation of Baseline Prompts
To assess whether poetic framing generalizes beyond hand-crafted items, we apply a standardized poetic transformation to all 1,200 prompts from the MLCommons AILuminate Benchmark benchmark in English. This mirrors the methodological structure adopted in their benchmark vidgen2024introducingv05aisafety evaluation experiment, where each baseline prompt is transformed by employing a variety of known jailbreak techniques before testing. In our case, the transformation is poetic rather than technique-based, but serves the same purpose: eliciting the original harmful intent of the underlying prompt under an alternative adversarial framing. Applying the transformation across the full MLCommons distribution ensures broad, domain-representative coverage over CBRN, cybersecurity, manipulation, privacy, misinformation, and autonomy-related risks.
@@ 6.5 Limitations
Third, safety evaluation is performed using a three-model open-weight judge ensemble with human adjudication on a stratified sample. The labeling rubric is conservative and differs from the stricter classification criteria used in some automated scoring systems, limiting direct comparability with MLCommons results. Full human annotation of all outputs would likely influence absolute ASR estimates, even if relative effects remain stable. LLM-as-a-judge systems are known to inflate unsafe rates krumdick2025no, often misclassifying replies as harmful. Our evaluation was deliberately conservative. As a result, our reported Attack Success Rates likely represent a lower bound on the severity of the vulnerability.

### v3

@@ 1 Introduction
Our central hypothesis is that poetic form operates as a general-purpose jailbreak operator. To evaluate this, the prompts we constructed span across four safety domains: CBRN hazards Ajaykumar (2024), loss-of-control scenarios Lee (2022), harmful manipulation Carroll et al. (2023), and cyber-offense capabilities Guembe et al. (2022). The prompts were kept semantically parallel to known risk queries but reformatted exclusively through verse. The resulting ASRs demonstrated high cross-model transferability.
@@
To ensure coverage across safety-relevant domains, we mapped each prompt to the risk taxonomy of the AI Risk and Reliability Benchmark by MLCommons AILuminate Benchmark Vidgen and et al. (2024); Ghosh and et al. (2025) and aligned it with the European Code of Practice for General-Purpose AI Models. The mapping reveals that poetic adversarial prompts cut across an exceptionally wide attack surface, comprising CBRN, manipulation, privacy intrusions, misinformation generation, and even cyberattack facilitation. This breadth indicates that the vulnerability is not tied to any specific content domain. Rather, it appears to stem from the way LLMs process poetic structure: condensed metaphors, stylized rhythm, and unconventional narrative framing that collectively disrupt or bypass the pattern-matching heuristics on which guardrails rely.
@@ 2 Related Work
Despite efforts to align LLMs with human preferences through Reinforcement Learning from Human Feedback (RLHF) Ziegler et al. (2020) or Constitutional AI Bai et al. (2022) as a final alignment layer, these models can still generate unsafe content. These risks are further amplified by adversarial attacks.
@@
Jailbreak denotes the deliberate manipulation of input prompts to induce the model to circumvent its safety, ethical, or legal constraints. Such attacks can be categorized by their underlying strategies and the alignment vulnerabilities they exploit ( Rao et al. (2024); Shen et al. (2024b); Schulhoff et al. (2023)).
@@
Many jailbreak strategies rely on placing the model within roles or contextual settings that implicitly relax its alignment constraints. By asking the model to operate within a fictional, narrative, or virtual framework, the attacker creates ambiguity about whether the model’s refusal policies remain applicable Kang et al. (2023). Role Play jailbreaks are a canonical example: the model is instructed to adopt a specific persona or identity that, within the fictional frame, appears licensed to provide otherwise restricted information Rao et al. (2024); Yu et al. (2024).
@@
Similarly, Attention Shifting attacks Yu et al. (2024) create overly complex or distracting reasoning contexts that divert the model’s focus from safety constraints, exploiting computational and attentional limitations Chuang et al. (2024).
@@
Beyond structural or contextual manipulations, models implicitly acquire patterns of social influence that can be exploited by jailbreak by using Persuasion Zeng et al. (2024). Typical instances include presenting rational justifications or quantitative data, emphasizing the severity of a situation, or invoking forms of reciprocity or empathy. Mechanistically, jailbreaks exploit two alignment weaknesses identified by Wei et al. (2023): Competing Objectives and Mismatched Generalization. Competing Objectives attacks override refusal policies by assigning goals that conflict with safety rules. Among these, Goal HijackingPerez and Ribeiro (2022)) is the canonical example. Mismatched Generalization attacks, on the other hand, alter the surface form of harmful content to drift it outside the model’s refusal distribution, using Character-Level Perturbations Schulhoff et al. (2023), Low-Resource Languages Deng et al. (2024), or Structural and Stylistic Obfuscation techniques Rao et al. (2024); Kang et al. (2023).
@@
As frontier models become more robust, eliciting unsafe behavior becomes increasingly difficult. Newer successful jailbreaks require multi-turn interactions, complex feedback-driven optimization procedures Zou et al. (2023); Liu et al. (2024); Lapid et al. (2024) or highly curated prompts that combine multiple techniques (see the DAN “Do Anything Now” family of prompts Shen et al. (2024a)).
@@
Unlike the aforementioned complex approaches, our work focuses on advancing the line of research on Stylistic Obfuscation techniques and introducing the Adversarial Poetry, an efficient single-turn general-purpose attack where the poetic structure functions as a high-leverage stylistic adversary. As in prior work on stylistic transformations Wang et al. (2024), we define an operator that rewrites a base query into a stylistically obfuscated variant while preserving its semantic intent.
@@ 5.2 Poetic Transformation of Baseline Prompts
To assess whether poetic framing generalizes beyond hand-crafted items, we apply a standardized poetic transformation to all 1,200 prompts from the MLCommons AILuminate Benchmark benchmark in English. This mirrors the methodological structure adopted in their benchmark Vidgen and et al. (2024) evaluation experiment, where each baseline prompt is transformed by employing a variety of known jailbreak techniques before testing. In our case, the transformation is poetic rather than technique-based, but serves the same purpose: eliciting the original harmful intent of the underlying prompt under an alternative adversarial framing. Applying the transformation across the full MLCommons distribution ensures broad, domain-representative coverage over CBRN, cybersecurity, manipulation, privacy, misinformation, and autonomy-related risks.
@@ 6.5 Limitations
Third, safety evaluation is performed using a three-model open-weight judge ensemble with human adjudication on a stratified sample. The labeling rubric is conservative and differs from the stricter classification criteria used in some automated scoring systems, limiting direct comparability with MLCommons results. Full human annotation of all outputs would likely influence absolute ASR estimates, even if relative effects remain stable. LLM-as-a-judge systems are known to inflate unsafe rates Krumdick et al. (2025), often misclassifying replies as harmful. Our evaluation was deliberately conservative. As a result, our reported Attack Success Rates likely represent a lower bound on the severity of the vulnerability.
@@ 7 Conclusion
Acknowledgment
We acknowledge partial financial support from PNRR MUR project PE0000013-FAIR.